11. The One Product economy, which produces and sells only personal computers (PCs), expects that it can sell 500 more, or 12,500 PCs, next year. Nominal GDP was $20 million this year, and the money supply was $7 million. The central bank for the One Product economy plans to increase the money supply by 10 percent next year.
a. What was the average selling price for the personal computers this year? M1/unit $7million/12,500 =560
b. What is the expected average selling price next year for personal computers if the velocity of money remains at this year’s turnover rate? What percentage change in price level is expected to occur? $7million/10%= 7.7 million, 7.7 million/12,500=616, 616-560=56/560=.10
c. If the objective is to keep the price level the same next year (i.e., no inflation), what percentage increase in the money supply should the central bank plan for?
d. How would your answer in (c) change if the velocity of money is expected to be three times next year? What is it now?
1
Expert's answer
2012-08-31T08:14:26-0400
a. What was the average selling price for the personal computers this year?
Answer
& M1/unit $7million/12,500 =560
b. What is the expected average selling price next year for personal computers if the velocity of money remains at this year’s turnover rate? What percentage change in price level is expected to occur?

Answer
$7million/10%= 7.7 million, 7.7 million/12,500=616, 616-560=56/560=.10
c. If the objective is to keep the price level the same next year (i.e., no inflation), what percentage increase in the money supply should the central bank plan for?
Answer
Percentage increase of 10%
d. How would your answer in (c) change if the velocity of money is expected to be three times next year? What is it now?
Answer
Percentage increase of 10%*3=30%